OpenStudy (anonymous):

Riley noticed an aquarium in his father’s office. The aquarium had the dimensions 16 in. by 8.5 in. by 10.5 in. The formula for volume is: V = l × w × h. Another aquarium in the building has dimensions that are each triple the size of the aquarium in Riley’s father’s office. Riley thought that the volume would also triple. Is Riley correct? How many times greater is the volume of the larger aquarium than the smaller? Show your work and explain your reasoning.

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

you should do this in math

OpenStudy (anonymous):

I believe that riley is correct.

OpenStudy (anonymous):

@PyroYolka I Concur.

OpenStudy (kamibug):

Riley is incorrect. The aquarium is 27 times bigger. Multiply each original dimension by 3 and that'll give you the new dimensions. Use those new dimensions to solve for the new volume. Then divide the original volume by the new volume to see how much bigger it is! :D 16*3=48 8.5*3=25.5 10.5*3=31.5 48*25.5*31.5=38556 Now divide the new volume by the original volume... 38556/1428 = 27 So the bigger aquarium is 27 times bigger than the smaller one! :)